OpenStudy (anonymous):

If a sprinkler waters 1 over 15 of a lawn in 1 over 5 of an hour, how much time will it take to water the entire lawn? 2 hours 3 hours 10 hours 15 hours

hero (hero):

@_greatmath7 giving answers is against CoC

OpenStudy (anonymous):

I never saw the answer, yay :)

OpenStudy (anonymous):

@Hero can you walk me through it?

hero (hero):

Well, do you know how to figure out how much 1/5 of an hour is in minutes?

OpenStudy (anonymous):

not really

hero (hero):

Well let's do it this way: 1/5 of an hour = 1/5 times 60 minutes = 60/5 minutes.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

60/5 would be 13 right?

OpenStudy (anonymous):

no no 12?

hero (hero):

Yes, correct, now think of it this way: Every 12 minutes, 1/15 of the lawn is watered so... The first 12 minutes, 1/15 is watered After 24 minutes, 2/15 is watered ... After 36 minutes 3/15 is watered ... So as you can see, once we get to 15/15, the whole lawn will be watered which will be a total of 15 intervals until the whole lawn is watered. The shortcut to calculate this is to simply take 12 and multiply it by 15. So 12×15 minutes. Then divide that amount by 60 to convert back to hours.

hero (hero):

Sorry, had to correct a typo.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so then it would be 15 right?

hero (hero):

What kinds of calculations have you made so far? If you did what I explained, there's no way you get 15. Show me what you did to get 15.

OpenStudy (anonymous):

well i didnt get 15 i got 30. i thought that would be 3 hours but 15 made more sense. this is what is did..

hero (hero):

Make sure you read the instructions I wrote exactly first.

hero (hero):

15 is just the number of intervals it takes to water the entire lawn.

OpenStudy (anonymous):

12 x 15 = 180 180/60 = 3

hero (hero):

Yes, that's better.

OpenStudy (anonymous):

oh i said 30 before >_<

OpenStudy (anonymous):

i meant 3. Would that be 3 hours then?

hero (hero):

Yes, it would.

OpenStudy (anonymous):

ok thank you i will do what i did for this on other questions. i have so many like this

hero (hero):

Good luck with it.

OpenStudy (anonymous):

Thank you for the help. :)